OpenStudy (anonymous):

Which of the following equations could be the result of using the comparison method to solve the system shown? x + 2y = 6 x - 4y = 8 2y - 6 = 4y + 8 6 - 2y = -4y + 8 6 - 2y = 4y + 8

OpenStudy (anonymous):

i'd say the top one

OpenStudy (anonymous):

To do this problem you need to solve for x in the top equation (x + 2y = 6). Then, plug in what you get for x in the top equation into the bottom equation. You know you need to solve for X and plug it into the 2nd equation because your options to choose from (the bottom three equations) only have Ys in them. I.e., you need to get rid of X. Then, pick the one equation from the bottom three that represents what you just did with the system of equations. For example:

OpenStudy (anonymous):

silly method though

OpenStudy (anonymous):

x + 2y = 6 Solve for x: x = (6 - 2y) Plug into 2nd equation: (6 - 2y) - 4y = 8 Which is equivalent to: 6 - 2y = 4y + 8

OpenStudy (anonymous):

than you @test203957

OpenStudy (anonymous):

isn't that called substitution?

OpenStudy (anonymous):

@zimmah yes, we are substituting (6-2y) for x and replacing it in the 2nd equation. Good :)

OpenStudy (anonymous):

@torchriswil NP!

OpenStudy (anonymous):

his question was comparison method or whatever it is.

OpenStudy (anonymous):

Oh! I don't even remember what that is. Let me look it up.

OpenStudy (anonymous):

comparison method and substitution method are similar yes but there are differences

OpenStudy (anonymous):

i prefer substitution myself as well, but there's another method as well. Either way the question is asked is in an overly confusing method (at least for me it seems more confusing than it needs to be)

OpenStudy (anonymous):

Thanks for bringing that up. I'll look it up as a review because I'm curious to know. I can tell you that whatever it is, I either use it intuitively without knowing it, or never use it at all. And I'm an engineering major.

OpenStudy (anonymous):

i guess the method they use is this: x + 2y = 6 x - 4y = 8 x = -2y +6 x = 4y+8 -2y+6=4y+8

OpenStudy (anonymous):

Yeah that's definitely it. I guess you solve for x in both equations and then set them equal to each other!

OpenStudy (anonymous):

i also study engineering, aviation engineering to be precise

OpenStudy (anonymous):

Cool!

OpenStudy (anonymous):

thank you guys both of you

OpenStudy (anonymous):

no problem